# Unveiling the Hubble Constant through Galaxy Cluster Gas Mass Fractions

### v1

@@ III Results
We conduct a Bayesian statistical analysis to constrain $H_{0}$ using the Markov Chain Monte Carlo (MCMC) method. We employ the emcee sampler Foreman-Mackey et al. (2013) to generate the chains and use the GetDist Python package Lewis (2019) for chain processing.
By defining the vector $\bm{\Delta}\equiv H_{0}(1+\bm{z}^{*})^{2}[\bm{D}_{A}(\bm{z}^{*})]^{\rm GC}-[\bm{H}_{0}\bm{D}_{L}(\bm{z}^{*})]^{\rm SN}$, the $\chi^{2}$ function can be written as
@@
$(\sigma_{D_{A},i})^{2}=\left(\frac{2}{3}\frac{H_{0}(1+z_{i})^{2}[D_{A}(z_{i})]^{\rm GC}\sigma_{f_{\rm gas},i}}{f_{{\rm gas},i}}\right)^{2}.$ (17)
Finally, we express the likelihood function of our analysis as
@@
where ${\bm{\Theta}}=\{H_{0},\Omega_{b},\Omega_{m},K_{0},K_{1},\gamma_{0},\gamma_{1}\}$ is the vector of the free parameters used in the analysis. Note that the exact set of parameters depends on the gas mass fraction sample and on the $\Omega_{b}/\Omega_{m}$ prior.
@@
$\omega_{b,0}$ $\mathcal{N}[0.02237;0.0015]$ CMB
$\omega_{m,0}$ $\mathcal{N}[0.1200;0.0012]$ CMB
@@
The results are as follows:
  • GMF1: Utilizing the galaxy clustering prior yields $H_{0}=71.4^{+10.5}_{-8.9}$ km/s/Mpc, while employing the Planck prior results in $H_{0}=68.1^{+7.7}_{-7.0}$ km/s/Mpc.
  • GMF2: With the galaxy clustering prior, we find $H_{0}=75.5^{+9.9}_{-7.9}$ km/s/Mpc, and with the Planck prior, $H_{0}=72.8^{+5.2}_{-5.0}$ km/s/Mpc.
@@ III.1 Estimating $H_{0}$ using simulated data
Refer to captionRefer to caption
Figure 5: Left Panel: Redshift distribution of the GMF1-like simulated data. Right Panel: Precision in $H_{0}$ [km/s/Mpc] as a function of the number of $f_{\text{gas}}$ observations. Each line represent a specific value of the average relative error of the simulated data.
@@
  • As expected that the gas mass fraction in galaxy cluster be a constant, in addition to the relative error, we also consider the dispersion of the real data around its mean value along the redshift.
Taking into account the characteristics of the real data, we perform 100 simulations with 40, 80, 160, 320, and 500 $f_{\text{gas}}$ data points. Given that the average relative error in the real data is approximately 12%, we investigate the impact of improved precision in $f_{\text{gas}}$ data by considering simulated data sets with average relative errors of 12%, 6%, and 3%, while maintaining their fitted redshift evolution. The results of this analysis are presented in the right panel of Fig. 5. Assuming that future measurements maintain the same average relative error as current samples, we can project that, with an optimistic scenario of 500 $f_{\text{gas}}$ measurements, an accuracy of 5% in $H_{0}$ could be achieved. Furthermore, if future precision improves by factors of 2 and 3 over current levels, $H_{0}$ could be measured with accuracies of 4.3% and 3.5%, respectively. A similar conclusion is found for precision in terms of GMF2-like samples. Therefore, for consistency in presentation, we have retained this analysis for our main sample, the GMF1-like data samples. Importantly, in an optimistic observational scenario, we may achieve a competitive $H_{0}$ accuracy in the near future.
@@ IV Final Remarks
Our main analysis yields the following Hubble constant estimates: $H_{0}=71.4^{+10.5}_{-8.9}$ km/s/Mpc for GMF1 and $H_{0}=75.5^{+9.9}_{-7.9}$ km/s/Mpc for GMF2, both at 68% CL (see Fig. 4). Due to their large uncertainties, these constraints are consistent with the measurements provided by the SH0ES team and the Planck satellite, indicating potential agreement despite the inherent errors in galaxy cluster observations. It is widely recognized that a deeper understanding of intra-cluster gas and dark matter properties is essential to enhance the utility of galaxy clusters as cosmological probes.
Finally, the impact on the $H_{0}$ determination via data Monte Carlo simulation was explored. Our simulations showed that future gas mass fraction measurements could achieve a precision of up to 3.5% for $H_{0}$ (see Fig. 5). Therefore, as statistical and systematic uncertainties in galaxy cluster observations diminish, the methodology employed here, leveraging gas mass fraction measurements, holds promise for improving constraints on the Hubble constant in a minimal model-independent manner.
@@ Acknowledgements.
RFLH thanks the financial support from the Conselho Nacional de Desenvolvimento Científico e Tecnologico (CNPq, National Council for Scientific and Technological Development) under the project No. 309132/2020-7. RCN thanks the financial support from the CNPq under project No. 304306/2022-3, and the Fundação de Amparo à Pesquisa do Estado do RS (FAPERGS, Research Support Foundation of the State of RS) for partial financial support under project No. 23/2551-0000848-3.

### v2

@@ III Results
We conduct a joint Bayesian statistical analysis of the two aforementioned gas mass fraction data samples and the Markov Chain Monte Carlo (MCMC) method to constrain $H_{0}$ with all data available. We employ the emcee sampler Foreman-Mackey et al. (2013) to generate the chains and use the GetDist Python package Lewis (2019) for chain processing. In order to avoid a double counting in the statistical analysis, we exclude the common clusters from the GMF2 sample: Abell2029, Abell1835, Abell2204, J1415.1+3612, RXCJ2129.6+0005, RXJ1347.5-1145 and SPT-CLJ2043-5035.
By defining the vector $\bm{\Delta}\equiv H_{0}(1+\bm{z}^{*})^{2}[\bm{D}_{A}(\bm{z}^{*})]^{\text{GC}}-[\bm{H}_{0}\bm{D}_{L}(\bm{z}^{*})]^{\text{SN}}$, the $\chi^{2}$ function can be written as
@@
$(\sigma_{D_{A},i})^{2}=\left(\frac{2}{3}\frac{H_{0}(1+z_{i})^{2}[D_{A}(z_{i})]^{\text{GC}}\sigma_{f_{\text{gas}},i}}{f_{{\text{gas}},i}}\right)^{2}.$ (18)
Finally, we express the joint likelihood function of our analysis as
@@
where ${\bm{\Theta}}=\{H_{0},\omega_{b}/\omega_{m},K_{0}^{r_{2500}},K_{1}^{r_{2500}},\gamma_{0}^{r_{2500}},\gamma_{1}^{r_{2500}},\alpha,$ $K_{0}^{r_{500}},\gamma_{0}^{r_{500}}\}$ is the vector of the free parameters used in the analysis. It is worth to mention that we treat the gas mass calibration parameter, $K$, and the depletion factor parameter, $\gamma$, independently because each one corresponds to models with different spherical radii.
@@
$\omega_{b,0}/\omega_{m,0}$ $\mathcal{N}[0.1564,0.001596]$ CMB
@@
The results of the joint analysis of the GMF1 and GMF2 are as follows:
  • Utilizing the galaxy clustering prior yields $H_{0}=72.7^{+6.3}_{-5.6}$ km/s/Mpc.
  • Utilizing the Planck prior, $H_{0}=79.1^{+4.3}_{-3.6}$ km/s/Mpc.
@@
The authors of the Ref.Mantz et al. (2021) considered the six clusters with $z<0.16$ of the GMF1 sample, $K_{0}=0.93\pm 0.11$, a Gaussian prior on $\Omega_{b}/\Omega_{m}$ from the Planck results (considering the flat $\Lambda$CDM model) and marginalized over a non-flat cosmological model with $\omega$ free (dark energy equation parameter). It was estimated $H_{0}=72.22\pm 6.7$ km/s/Mpc. As one may see, their estimate is in full agreement with ours. However, no specific cosmological framework was considered in our estimates, we marginalized on a possible evolution of the astrophysical parameters or mass dependence ($\gamma(z)$ $K(z)$ and $\alpha$) and considering all galaxy clusters in a joint statistical analysis of the GMF1 and GMF2 samples.
@@ III.1 Estimating $H_{0}$ using simulated data
Figure 5: Left Panel: Redshift distribution of the GMF1-like simulated data. Right Panel: Precision in $H_{0}$ as a function of the number of $f_{\text{gas}}$ observations. Each line represents a specific value of the average relative error of the simulated data.
@@
  • As the gas mass fraction in the galaxy cluster is expected to be constant, in addition to the relative error, we also consider the dispersion of the actual data around its mean value across different redshifts.
Taking into account the characteristics of the real data, we perform 100 simulations with 40, 80, 160, 320, and 500 $f_{\text{gas}}$ data points. Given that the average relative error in the real data is approximately 12%, we investigate the impact on $H_{0}$ estimates of improved precision in $f_{\text{gas}}$ data by considering simulated data sets with average relative errors of 12%, 6%, and 3%, while maintaining their fitted redshift evolution and astrophysical and cosmological parameters unaltered. The results of $H_{0}$ simmetrized errors Schmelling (2000) performing this analysis are presented in the right panel of Fig. 5. Assuming that future measurements maintain the same average relative error as current samples, we can project that, with an optimistic scenario of 500 $f_{\text{gas}}$ measurements, an accuracy of 7.3% in $H_{0}$ could be achieved. Furthermore, if future $f_{\text{gas}}$ data precision improves by factors of 2 and 4 over current levels, $H_{0}$ could be measured with precision of 6.2% and 5.0%, respectively. It is worth mentioning that in this analysis, the precision of individual $H_{0}$ estimates obtained from each $f_{\text{gas}}$ data point is limited by the precision of the astrophysical and cosmological parameters. However, by combining all $H_{0}$ information statistically, the final estimate may achieve higher precisions. A similar conclusion is found for precision in terms of GMF2-like samples. Therefore, for consistency in presentation, we have retained this analysis for our main sample, the GMF1-like data samples. Importantly, in an optimistic observational scenario, we may achieve a competitive $H_{0}$ accuracy in the near future.
@@ IV Final Remarks
Our analysis yields the following Hubble constant estimate: $H_{0}=72.7^{+6.3}_{-5.6}$ km/s/Mpc for the compilation of the samples GMF1 and GMF2 at 68% CL (see Fig. 4). Due to their large uncertainties, these constraints are consistent with the measurements provided by the SH0ES team and the Planck satellite, indicating potential agreement despite the inherent errors in galaxy cluster observations. It is widely recognized that a deeper understanding of intra-cluster gas and dark matter properties is essential to enhance the utility of galaxy clusters as cosmological probes. In general, the mass of a galaxy cluster obtained using strong gravitational lensing tends to be greater than the mass derived from the hypothesis of hydrostatic equilibrium. In other words, the hydrostatic equilibrium hypothesis adopted in these studies may be underestimating the gas mass fraction measurements and, consequently, overestimating the Hubble constant estimates. However, the galaxy clusters considered in analyses are limited to the most dynamically relaxed, massive clusters known. In turn, this restriction is critical for reducing systematic from hidrostatical equilibrium and spherical symmetry.
Finally, the impact on the $H_{0}$ determination via data Monte Carlo simulation was explored. Our simulations showed that future gas mass fraction measurements could achieve a precision of up to 5% for $H_{0}$ (see Fig. 5). Therefore, as statistical and systematic uncertainties in galaxy cluster observations diminish, the methodology employed here, leveraging gas mass fraction measurements, holds promise for improving constraints on the Hubble constant in a minimal model-independent manner.
@@ Acknowledgements.
The authors thank the referee and Susana Landau for valuable comments about this work. RFLH thanks the financial support from the Conselho Nacional de Desenvolvimento Científico e Tecnologico (CNPq, National Council for Scientific and Technological Development) under the project No. 309132/2020-7. RCN thanks the financial support from the CNPq under project No. 304306/2022-3, and the Fundação de Amparo à Pesquisa do Estado do RS (FAPERGS, Research Support Foundation of the State of RS) for partial financial support under project No. 23/2551-0000848-3. LRC also thanks to CNPQ support under the project No. 169625/2023-0. ..